can you help me solve this problem: 13-8v = 5v+3
Answer by tjnw79(57)

13-8v = 5v+3
first subtract 3 on both sides
-8v+(13)-3=5v+3-(3) to get
10-8v=5v
now add 8v to both sides
10-8v+(8v)=5v+(8v)
10=13v
becuse 13 is times v we need to divde 13 on both sides to get you answer
10/13=v